Find the mean square radius of 56Ni nucleus assuming it to be a spherical nucleus with sharp boundary


1
Expert's answer
2021-09-09T10:52:05-0400

Given:

"A=56"

"r_0=1.3*10^{-15}\:\rm m"


The radius of nucleus

"R=r_0A^{1\/3}"

"R=1.3*10^{-15}{\:\rm m}*56^{1/3}=5.0*10^{-15}{\:\rm m}\\=5.0\:\rm fm"